# OntoPortal-Astro, a Semantic Artefact Catalogue for Astronomy

### v1

@@ 3.1 Exoplanets
Currently, the uat and ivoauat thesauri are published as SKOS vocabularies. The objtyp and tarcls vocabularies are draft SAs prepared respectively by the IVOA Semantics working group, and by the Solar System interest group of the IVOA with the Europlanet team, and are being developed as OWL ontologies. Mapping these SAs is not a difficult task, but collecting them into a single OntoPortal instance makes is easier.
@@ 3.5 Scientific Knowledge Graphs
Connecting research outputs to persons, institutions and funders is a key task for enabling the FAIR principles, discovering research products or measuring the scientific impact of the projects and infrastructures. Several tools are implementing so called SKG, which can mean either Scholarly Knowledge Graph (such as with the OpenAIRE Graph infrastructure [^19] https://graph.openaire.eu, or Scientific Knowledge Graph (such as with the SciLake project [^20] https://scilake.eu). Building efficient and FAIR knowledge graphs requires to use well-defined and cross-domain ontologies, so that term URIs can be used as search keys in queries (see, e.g., the “exoplanet” use case). OntoPortal implements a very rich metadata model to describe SAs — based on the MOD metadata vocabulary [^21] https://github.com/FAIR-IMPACT/MOD (Dutta et al., 2017) — which will support the formal and semantic description of astronomy SAs in SKGs. Plus OntoPortal embeds a FAIRness assessment tool based on the O’FAIRe methodology (Ontology FAIRness Evaluation, Amdouni et al., 2022) to help in the selection of SAs.
@@ 3.6 Models and observations
Astronomy is an observation-based science, since the environmental conditions of the observed objects and phenomena can’t be controlled. It is thus necessary to try and model the possible environmental conditions leading to the observations. Within the IVOA, the celestial astronomy community has developed the Simulation Data Model (Lemson et al., 2012), which can describe the astrophysical simulation parameters. However, due to its complexity, it has not been adopted by many teams. In the heliophysics community, the need to coupling measurements to simulations has been identified as a key issue (Masson et al., 2024), either to discover and reuse observational data as input for environment models, or to compare observations to output parameters of those models. Hence, the observed and modelled physical parameters must be described the same way. The same reasoning applies to planetary environment, such as the global atmospheric models of planets, or the planetary magnetosphere models (see, e.g., Erard et al., 2018; Achilleos et al., 2019; Azria et al., 2021; Trompet et al., 2022). The topic of representing observable properties (or scientific variables) is very present in multiple OntoPortal community such as Earth sciences (EarthPortal), Agri-food (AgroPortal) or Ecology (EcoPortal). In a near future, OntoPortal technology will support standard ways to represent variables (such as the Extensible Observation Ontology (OBOE, Madin et al., 2007) or the InteroperAble Descriptions of Observable Property Terminology [^22] https://i-adopt.github.io (I-ADOPT)).
@@ 4.1 Review of pre-exisiting Semantic Artefact catalogues
A recent landscape analysis of catalogues containing SA has identified three among the astronomy community: the IVOA vocabularies, the NASA/PDS data standards and the ESPAS Vocabulary broker (Jonquet and Grau, 2024b). SA catalogues have been classified according to their typology and the level of service offered. Listing SAC correspond to a simple online listing of SA, while Library represent structured online listing and the Repository is considered as an advance web application (Jonquet and Grau, 2024a). The IVOA vocabulary [^23] https://ivoa.net/rdf/ web page is considered as a SA library, which gathers the vocabularies developed and maintained by the IVOA Semantics working group. The NASA/PDS data standard [^24] https://pds.nasa.gov/datastandards web page is a SA listing, which provides all version of the NASA/PDS information model, and the associated vocabularies. The ESPAS Vocabulary Broker, which is consider as retired (Jonquet and Grau, 2024a), and is not accessible anymore at the time of writing, is a SA repository. It was providing access to the terms from the SPASE schema, the ESPAS project ontology, the GEMET ontology and the UAT thesaurus. Moreover, the ESPAS vocabulary used to offer mapping between terms of ontologies among the repository. Among those, solely the IVOA vocabulary library partly meets the SA catalogues FAIRness criteria proposed in Jonquet and Grau (2024a). Hence, no pre-existing solution satisfies the needs for a common astronomy SA catalogue. There is thus a need of harmonising and synchronising the different initiatives to avoid technical and semantic silos.
@@ 4.2 OntoPortal-Astro, an astronomy community OntoPortal
http://ontoportal-astro.eu [^25] At the time of writing, this server presents the astronomy OntoPortal prototype developed during the FAIR-IMPACT project, but will be replaced by the OntoPortal-Astro within a few months.
@@ 4.3 Improvement of IVOA vocabularies
Importing the IVOA vocabularies into our astronomy OntoPortal instance revealed minor design issues in the IVOA vocabularies management and their associated RDF artefacts, as defined in Demleitner et al. (2023). We fixed the SKOS vocabularies, which were missing top level concepts. This prevented tools to correctly build the concept hierarchies out of the ontology. We also fixed the way the URI were referenced internally in the ontologies. Finally, we started to add more metadata to our vocabularies, such as, preferred namespace prefix and URIs, using the multiple properties harvested and promoted by MOD. We also initiated a mapping exercise between some of our properties with more commonly used ones. For instance, we have two properties used to handle deprecated terms and redirect to a replacement term: ivoasem:deprecated and ivoasem:useInstead, as part of our IVOA Semantics vocabulary (IVOASEM). Section B shows a possible implementation of these mappings.
@@ 4.4 Strong heliophysics support
Thanks to the Transform to Open Science (TOPS) project, the NASA contributors to IHDEA started several projects aiming at improving semantic interoperability of the heliophysics community. A first example is the conversion of the SPASE XML Schema into an OWL ontology [^26] https://github.com/polyneme/topst-spase-rdf-tools. A second ongoing work is the mapping of the SPASE XML Schema to schema.org [^27] https://github.com/clnsmth/soso, in coordination with Science on Schema.org (Shepherd et al., 2024).
@@ 5 Challenges and impact
Figure 1: Impact factors and metrics for assessing the impact of the OntoPortal technology on the astronomy community. Figure extracted from Aubin et al. (2025).
@@ 6 Discussion
The development of the astronomy OntoPortal prototype has been truly interdisciplinary. It was enabled by the FAIR-IMPACT project, which has been funded European Open Science Cloud (EOSC) infrastructure programme. One of the goals of FAIR-IMPACT was to develop a multi-disciplinary federation of SA catalogues. The project selected the OntoPortal technology. Although the astronomy communities are mature communities, with well-established data ecosystems and solid bottom-up metadata governance, their use of web-semantics technologies is not so much developed, compared to the practices in place in other communities like environmental or life sciences. The collaboration between the astronomy community on the one hand, and INRAE and the OntoPortal alliance on the other hand, brought astronomy the missing pieces to fast-forward to the modern age of web-semantics, with state-of-art technologies. The collaborations initiated during this project will be continued, with upcoming activities on OntoPortal-Astro and other projects of the EOSC programme.
@@
As shown in Molinaro et al. (2020), in the frame of the ESCAPE project, the IVOA Registry has been connected to B2Find (a service provided by EUDAT, in the frame of EOSC): https://b2find.eudat.eu/organization/ivoa. This provides a new venue for finding IVOA services and collections. However, the current connection is not implementing any of the implicit IVOA semantics (like, e.g., the IVOA UAT keywords, which are available as pure text terms, and not URIs). The goal of the FAIR-IMPACT project was to expand FAIR Solutions across EOSC, including web-semantics technologies. Once OntoPortal-Astro is operational, further work with the IVOA and EUDAT/B2Find will have to be conducted to connect this new SA catalogue to the B2Find search portal.
@@
The developments around the astronomy OntoPortal prototype have been regularly discussed within the IVOA Semantics working group. While the working members are fully aware and positive about the open science challenges and the need for improving the astronomy SAs, the common ground is that the IVOA vocabularies are tailored for enabling interoperability within the IVOA. Hence, any development on the IVOA vocabularies shall be made backward compatible with the current uses and practices and shall not impose extra burden on the maintainers, tool developers and users of the vocabularies.
@@
The IVOA vocabularies are setup so that they have as little external dependencies as much as possible. The main example is the UAT, which is developed independently from the IVOA. The Unified Astronomy Thesaurus (IVOA rendering) is a regularly updated copy of the UAT thesaurus. The IVOA version includes the relations to the original UAT terms, so that linked data applications can connect resources using the same terms. However, in the unfortunate event of discontinuation of the UAT reference thesaurus, the IVOA interfaces based on those terms will continue to work. The current design is thus trying to keep a strong infrastructure, as well as allowing linked and open data applications. For this, OntoPortal will also be a long term solution in case a vocabulary cease to be supported (e.g., URIs are not resolvable anymore) as the technology include a URI resolution and negotiation service for each hosted SA.
@@
Improving the FAIRness of the astronomy vocabulary is also one of the key outcome of this study. We used FAIR assessment tools dedicated to ontologies and SAs, like FOOPS! [^28] https://foops.linkeddata.es/FAIR_validator.html (Garijo et al., 2021). Among all the assessment criteria, we can list the following aspects. The term URIs should be persistent, but the tools only consider a limited set of http domains as persistent. The metadata should include a list of contributors, for the ontology and the individual terms. It should also include provenance metadata. A versioning schema should also be in place. A preferred prefix should be advertised. Some criteria can easily be set up, and others (like the persistent identifier one) would require coordination with the FAIR assessment tool maintainers. On this aspect, the future OntoPortal-Astro will include O’FAIRe tool which provides a nice reporting on how to improve the FAIRness of each SAs.
@@
This study is a first step towards improved management practices of SAs for astronomy. We focused here on vocabularies and ontologies, which are already in use by the astronomy communities. A further step is to ease and extend their usage for describing and annotating research objects. The Astronomy community pilot developed within the OSTrails project will make use of the FAIR SAs developed using OntoPortal for connecting resources through Science Knowledge Graph interfaces. Finally, another framework identified for annotated research products with description of measurements is I-ADOPT. It is proposing a simple template with design patterns and is making use of existing ontologies. We participated to the I-ADOPT Variable Modelling Challenge in 2024, with an example variable “electron density in the solar wind”. From this simple test, the framework seems applicable, and we will explore more examples with astronomy use cases.
@@
Keeping our astronomy ecosystem stable while opening up to linked data practices requires to build term mappings. Recently developed tools (see, e.g., MSCR – Metadata Schema and Crosswalk Registry, Kesäniemi et al., 2025) are promising resources for storing mappings, including more complex conditional schema mappings than the current capabilities of OntoPortal, as those developed during the SPASE XML Schema to Schema.org exercise. The mappings will then have to be stored, either in independent mapping registries (such as the MSCR or the native OntoPortal mapping repository) or integrated into our ontologies. Using artificial intelligence and large language models will also be tested for aligning vocabularies and building mappings.
@@
Since January 2025, the OSCARS cascading grant “Ontology Portal for Astronomy Linked-data” [^29] https://oscars-project.eu/projects/opal-ontology-portal-astronomy-linked-data (OPAL) started. Its goal is to build a real production service, OntoPortal-Astro for the astronomy community building on the preliminary results from FAIR-IMPACT. Through this project, we plan to gather, update and produce SAs for all the astronomy sub-communities. We will also explore how to better expose our SAs so that they can be reused. For this, OntoPortal-Astro’s team will join the Ontoportal Alliance, and the portal will federate with other OntoPortal instances, such as EarthPortal, which is serving ontologies we can relate our terms to (for the space weather community or the planetary surfaces). We also consider developing domain ontologies, which will allow us to annotate our research products in a more accurate manner.
@@ Acknowledgments
This work has been supported by the FAIR-IMPACT project, which received funding from the European Commission’s Horizon Europe Research and Innovation programme under grant agreement number 101057344. The authors acknowledge the Observatoire de Paris-PSL for hosting the OntoPortal prototype used for this work and the OntoPortal Alliance for their support and technology. The authors also thank Markus Demleitner (former chair of the IVOA Semantics working group), as well as the IVOA, IHDEA and IPDA groups for their helpful discussions and feedback. The authors also acknowledge support from Sophie Aubin from INRAE.

### v3

@@ 3.1 Exoplanets
Currently, the uat and ivoauat thesauri are published as SKOS vocabularies. The objtyp and tarcls vocabularies are draft SAs prepared respectively by the IVOA Semantics WG, and by the Solar System interest group of the IVOA with the Europlanet team, and are being developed as OWL ontologies, using either Protégé, or the IVOA vocabulary framework. Mapping these SAs is not a difficult task, but collecting them into a single OntoPortal instance makes it easier.
@@ 3.5 Scientific Knowledge Graphs
Connecting research outputs to persons, institutions and funders is a key task for enabling the FAIR principles, discovering research products or measuring the scientific impact of the projects and infrastructures. Several tools are implementing so called SKG, which can mean either Scholarly Knowledge Graph (such as with the OpenAIRE Graph infrastructure [^27] https://graph.openaire.eu), or Scientific Knowledge Graph (such as with the SciLake project [^28] https://scilake.eu). Moreover, the astronomy community registries, such as the IVOA registry or the SPASE registry, can also be named SKGs as they relate all the elements of the research ecosystem (data products, repositories, scholarly articles, institutions, persons, etc.)
Building efficient and FAIR knowledge graphs requires to use well-defined and cross-domain ontologies, so that term URIs can be used as search keys in queries (see, e.g., the “exoplanet” use case). OntoPortal-Astro will implements a very rich metadata model to describe SAs — based on the MOD metadata vocabulary [^29] https://github.com/FAIR-IMPACT/MOD (Dutta et al., 2017) — which will support the formal and semantic description of astronomy SAs in SKGs. Plus, it will embed a FAIRness assessment tool based on the O’FAIRe methodology (Ontology FAIRness Evaluation, Amdouni et al., 2022) to help in the selection of SAs. OntoPortal-Astro will thus facilitate the reuse of astronomy SAs making them more findable.
@@ 3.6 Models and observations
Astronomy is an observation-based science, since the environmental conditions of the observed objects and phenomena can’t be controlled. It is thus necessary to try and model the possible environmental conditions leading to the observations. Within the IVOA, the celestial astronomy community has developed the Simulation Data Model (Lemson et al., 2012), which can describe the astrophysical simulation parameters. However, due to its complexity, it has not been adopted by many teams. In the heliophysics community, the need to coupling measurements to simulations has been identified as a key issue (Masson et al., 2024), either to discover and reuse observational data as input for environment models, or to compare observations to output parameters of those models. Hence, the observed and modelled physical parameters must be described the same way. The same reasoning applies to planetary environment, such as the global atmospheric models of planets, or the planetary magnetosphere models (see, e.g., Erard et al., 2018; Achilleos et al., 2019; Azria et al., 2021; Trompet et al., 2022).
Describing observable properties (or scientific variables) within ontologies or other types of SAs, is a growing practice. It is present in multiple OntoPortal community instances such as Earth sciences (EarthPortal), Agri-food (AgroPortal) or Ecology (EcoPortal), using the Extensible Observation Ontology (OBOE, Madin et al., 2007) or the InteroperAble Descriptions of Observable Property Terminology [^30] https://i-adopt.github.io (I-ADOPT)). For instance, in EcoPortal, the ENVTHES (Environmental Thesaurus) already use the I-ADOPT representation to encode their scientific variable within these SAs  [^31] See, e.g., http://ecoportal.lifewatch.eu/ontologies/ENVTHES/30209.
@@ 4.1 Review of pre-exisiting Semantic Artefact Catalogues
A recent landscape analysis of catalogues containing SAs has identified three among the astronomy community: the IVOA vocabularies, the NASA/PDS data standards and the ESPAS Vocabulary broker (Jonquet and Grau, 2024b). SA catalogues have been classified according to their typology and the level of service offered. A listing corresponds to a simple online listing of SAs, while a library represents structured online listing and the a repository is considered as an advanced web application (Jonquet and Grau, 2024a). The IVOA vocabulary [^32] https://ivoa.net/rdf/ web page is considered as a SA library, which gathers the vocabularies developed and maintained by the IVOA Semantics WG. The NASA/PDS data standard [^33] https://pds.nasa.gov/datastandards web page is a SA listing, which provides all version of the NASA/PDS information model, and the associated vocabularies. The ESPAS Vocabulary Broker, which is consider as retired (Jonquet and Grau, 2024a), and is not accessible anymore at the time of writing, is a SA repository. It was providing access to the terms from the SPASE schema, the ESPAS project ontology, the GEMET ontology and the UAT thesaurus. Moreover, the ESPAS vocabulary used to offer mapping between terms of ontologies among the repository. Among those, solely the IVOA vocabulary library partly meets the SA catalogues FAIRness criteria proposed in Jonquet and Grau (2024a). Hence, no pre-existing solution satisfies the needs for a common astronomy SA catalogue. There is thus a need of harmonising and synchronising the different initiatives to avoid technical and semantic silos.
@@ 4.2 OntoPortal-Astro, an astronomy community OntoPortal
https://ontoportal-astro.eu [^34] At the time of writing, this server presents the astronomy OntoPortal prototype developed during the FAIR-IMPACT project, but will be replaced by the OntoPortal-Astro within a few months.
@@ 4.3 Improvement of IVOA vocabularies
Importing the IVOA vocabularies into our astronomy OntoPortal instance revealed minor design issues in the IVOA vocabularies management and their associated RDF artefacts, as defined in Demleitner et al. (2023). We fixed the SKOS vocabularies, which were missing top level concepts. This prevented tools to correctly build the concept hierarchies out of the ontology. We also fixed the way the URI were referenced internally in the ontologies. Finally, we started to add more metadata to our vocabularies, such as, preferred namespace prefix and URIs, using the multiple properties harvested and promoted by MOD. We also initiated a mapping exercise between some of our properties with more commonly used ones. For instance, we have two properties used to handle deprecated terms and redirect to a replacement term: ivoasem:deprecated and ivoasem:useInstead, as part of our IVOA Semantics vocabulary (IVOASEM).
@@ 4.4 Strong heliophysics support
Thanks to the Transform to Open Science (TOPS) project, the NASA contributors to IHDEA started several projects aiming at improving semantic interoperability of the heliophysics community. A first example is the conversion of the SPASE XML Schema into an OWL ontology [^35] https://github.com/polyneme/topst-spase-rdf-tools. A second ongoing work is the mapping of the SPASE XML Schema to schema.org [^36] https://github.com/clnsmth/soso, in coordination with Science on Schema.org (Shepherd et al., 2024; Ringuette et al., 2025). When implemented and operational, these mappings will require an easy access to SAs, based search interfaces on ontology terms, so that term URIs are seamlessly inserted in the metadata. OntoPortal-Astro will provide a search interfaces for each SA, including a programmatic access.
@@ 5 Challenges and impact
Figure 1: Impact factors and metrics for assessing the impact of the OntoPortal technology on the astronomy community. Figure adapted from Aubin et al. (2025).
@@ 6 Discussion
The development of the astronomy OntoPortal prototype has been truly interdisciplinary. It was enabled by the FAIR-IMPACT project, which has been funded European Open Science Cloud (EOSC) infrastructure programme. One of the goals of FAIR-IMPACT was to develop a multi-disciplinary federation of SA catalogues. The project selected the OntoPortal technology. Although the astronomy communities are mature communities, with well-established data ecosystems and solid bottom-up metadata governance, their use of semantic web technologies is not so much developed, compared to the practices in place in other communities like environmental or life sciences. The collaboration between the astronomy community on the one hand, and INRAE and the OntoPortal alliance on the other hand, brought astronomy the missing pieces to fast-forward to the modern age of web-semantics, with state-of-art technologies. The collaborations initiated during this project will be continued, with upcoming activities on OntoPortal-Astro and other projects of the EOSC programme.
@@
As shown in Molinaro et al. (2020), in the frame of the ESCAPE project, the IVOA Registry has been connected to B2Find (a service provided by EUDAT, in the frame of EOSC): https://b2find.eudat.eu/organization/ivoa. This provides a new venue [^37] In the sense of the SKG-IF “venue” definition: A publishing “gateway” used by an Agent to make their Research products available to others. for finding IVOA services and collections. However, the current connection is not implementing any of the implicit IVOA semantics (like, e.g., the IVOA UAT keywords, which are available as pure text terms, and not URIs). The goal of the FAIR-IMPACT project was to expand FAIR solutions across EOSC, including semantic web technologies. Once OntoPortal-Astro is operational, further work with the IVOA and EUDAT/B2Find will have to be conducted to connect this new SA catalogue to the B2Find search portal.
@@
The developments around the astronomy OntoPortal prototype have been regularly discussed within the IVOA Semantics WG. While the WG members are fully aware and positive about the open science challenges and the need for improving the astronomy SAs, the common ground is that the IVOA vocabularies are tailored for enabling interoperability within the IVOA. Hence, any development on the IVOA vocabularies shall be made backward compatible with the current uses and practices and shall not impose extra burden on the maintainers, tool developers and users of the vocabularies.
@@
The IVOA vocabularies are setup so that they have as little external dependencies as possible. The main example is the UAT, which is developed independently from the IVOA. The Unified Astronomy Thesaurus (IVOA rendering) is a regularly updated copy of the UAT thesaurus. The IVOA version includes the relations to the original UAT terms, so that linked data applications can connect resources using the same terms. However, in the unfortunate event of discontinuation of the UAT reference thesaurus, the IVOA interfaces based on those terms will continue to work. The current design is thus trying to keep a strong infrastructure, as well as allowing linked and open data applications. For this, OntoPortal-Astro will serve as a long term solution in case a vocabulary cease to be supported (e.g., URIs are not resolvable anymore), as the technology include a URI resolution and negotiation service for each hosted SA.
Even with the impulse of OntoPortal-Astro, the semantic artefact management may not change in all communities immediately. For instance the IVOA has already a well established SA governance, and the improvement should only be limited to the improvement of the SA metadata. For planetary sciences and heliophysics, the first step will be to include OWL or SKOS versions of their SAs, and then include this scheme into their SA governance. Currently, the heliophysics community is studying the mapping of their registry metadata records to schema.org, which requires to produce SAs and accessible concept URIs. In the planetary science community, the IPDA is currently not ready for such a shift, but the planetary surfaces OGC group should be able to produce SAs related to coordinate systems and projections.
@@
Improving the FAIRness of the astronomy vocabulary is also one of the key outcome of this study. We used FAIR assessment tools dedicated to ontologies and SAs, like FOOPS! [^38] https://foops.linkeddata.es/FAIR_validator.html (Garijo et al., 2021). Among all the assessment criteria, we can list the following aspects. The term URIs should be persistent, but the tools only consider a limited set of http domains as persistent. The metadata should include a list of contributors, for the ontology and the individual terms. It should also include provenance metadata. A versioning schema should also be in place, as discussed in Jonquet and Poveda-Villalón (2023). A preferred prefix should be advertised. Some criteria can easily be set up, and others (like the persistent identifier one) would require coordination with the FAIR assessment tool maintainers. On this aspect, the future OntoPortal-Astro will include O’FAIRe tool which provides a nice reporting on how to improve the FAIRness of each SAs.
@@
This study is a first step towards improved management practices of SAs for astronomy. We focused here on vocabularies and ontologies, which are already in use by the astronomy communities. A further step is to ease and extend their usage for describing and annotating research objects. The astronomy community pilot developed within the OSTrails project will make use of the FAIR SAs developed using OntoPortal for connecting resources through Science Knowledge Graph interfaces. Finally, another framework identified for annotated research products with description of measurements is I-ADOPT. It is proposing a simple template with design patterns and is making use of existing ontologies. We participated to the I-ADOPT Variable Modelling Challenge in 2024, with an example variable “electron density in the solar wind”. From this simple test, the framework seems applicable, but, a lot concepts related to astronomy are missing. We will explore more examples with astronomy use cases and identify which ontologies and concepts should be implemented to cover the needs.
@@
Keeping our astronomy ecosystem stable while opening up to linked data practices requires to build term mappings. Recently developed tools (see, e.g., MSCR – Metadata Schema and Crosswalk Registry, Kesäniemi et al., 2025) are promising resources for storing mappings, including more complex conditional schema mappings than the current capabilities of OntoPortal, as those developed during the SPASE XML Schema to Schema.org exercise. The mappings will then have to be stored, either in independent mapping registries (such as the MSCR or the native OntoPortal mapping repository) or integrated into our ontologies. The use of artificial intelligence and large language models will also be tested for aligning vocabularies and building mappings.
@@
Since January 2025, the OSCARS cascading grant “Ontology Portal for Astronomy Linked-data” [^39] https://oscars-project.eu/projects/opal-ontology-portal-astronomy-linked-data (OPAL) started. Its goal is to build a real production service, OntoPortal-Astro for the astronomy community building on the preliminary results from FAIR-IMPACT. Through this project, we plan to gather, update and produce SAs for all the astronomy sub-communities. We will also explore how to better expose our SAs so that they can be reused. The OntoPortal-Astro team will include ontologists and knowledge designers, and will be supported by an Advisory and User Group composed of experts from the various astronomy communities. OntoPortal-Astro will join the Ontoportal Alliance, and the portal will federate with other OntoPortal instances, such as EarthPortal, which is serving ontologies we can relate our terms to (for the space weather community or the planetary surfaces). We also consider developing domain ontologies, which will allow us to annotate our research products in a more accurate manner.
@@ Acknowledgments
This work has been supported by the FAIR-IMPACT project, which received funding from the European Commission’s Horizon Europe Research and Innovation programme under grant agreement number 101057344. The authors acknowledge the Observatoire de Paris-PSL for hosting the OntoPortal prototype used for this work and the OntoPortal Alliance for their support and technology. The authors also thank Markus Demleitner (former chair of the IVOA Semantics WG), as well as the IVOA, IHDEA and IPDA groups for their helpful discussions and feedback. The authors also acknowledge support from Sophie Aubin from INRAE.